A sample of helium has a temperature of 120. °C and pressure of 746 Torr. What is the temperature (in °C), when the pressure is 715 Torr?
1
Expert's answer
2019-10-25T05:31:03-0400

PV=nRT

P1/P2=T1/T2

T1=120+273.15=393.15K

746/715=393.15/T2

T2=376.81 K = 103.66 C

1P